# DeciMamba: Exploring the Length Extrapolation Potential of Mamba

### v2

@@ 1 Introduction
Figure 1: Improving Mamba Extrapolation with DeciMamba. We present a novel decimation mechanism tailored for Mamba. With our method Mamba is able to process sequences that are significantly longer than the ones trained on while enjoying reduced inference costs. (Left) Document Retrieval. (Right) Passkey Retrieval. In all plots, the purple line shows the training sequence length. Results are shown for DeciMamba-130M.
@@ 5 Experiments
Document Retrieval In this task the model receives a query and $N_{docs}$ randomly assorted documents, and its objective is to return the id of the document that holds the answer to the query. Our data is sampled from SQuAD v2 (rajpurkar2018squadv2). We train both Mamba-130m and DeciMamba-130m to retrieve out of $N_{docs}=11$ documents (about 2k tokens), and evaluate their performance with $N_{docs}\in[11,300]$ (2k to 60k tokens). The results are presented in Fig. 1 (left). While Mamba is able to retrieve until $N_{docs}=40$ documents, it suffers from a sharp decline when the amount of documents is further extended. On the contrary, DeciMamba is able to extrapolate to sequences that are more than x25 times longer than the ones seen during training ($\sim 250$ documents), and does so without any additional computational resources. See Sec. A.2 for more details.
Multi-Document QA. Next, we stay in a setting similar to Sec. 1 (Document Retrieval), yet increase the level of difficulty by asking the model to answer the query in free text (instead of retrieving the id of the most relevant document). We train Mamba-130m and DeciMamba-130m on the same dataset and present the F1 score between the generated response and the ground truth answers in Tab. 2. As in the previous task, when the amount of documents is close to the amount trained on, the models perform quite similarly, with a slight advantage towards DeciMamba. When the document amount significantly increases we can see that DeciMamba develops a clear advantage. Yet, this time, due to the task’s difficulty, its performance decays as well. Note that Mamba’s and DeciMamba’s performance is relatively modest compared to prior work due to the use of smaller models, which limits the presented text generation capabilities.
Table 2: Multi-Document QA. The table holds the F1 score between the ground truth annotations and the predictions of Mamba-130m and DeciMamba-130m (+Deci) across different number of documents. Both models were trained using 11 documents.
# Docs 5 10 20 40 60 80 100 120 140 160
Mamba 23.5 25.3 22.8 23.6 25.8 19.3 7.5 4.1 1.1 0.6
+Deci 25.6 26.2 24.4 25.3 22.74 19.1 18.1 12.9 10.0 5.0
@@ A.1 Passkey Retrieval
A.2 Document Retrieval
We train each model with data from SQuAD v2 (rajpurkar2018squadv2), which provides examples in the form of (Query, Document, Answer). Our training samples have the following form: $N_{docs}\times$$<$$Document$$>$; $<$$Answer$$>$, where $<$$Document$$>$ can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. $<$$Answer$$>$ holds the id of the golden document. In our setting $N_{docs}=12$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During Evaluation we use the same setting but vary the value of $N_{docs}$, between 12 and 300. We note that the length of an average document in SQuAD is a bit smaller than 200 tokens, so our average training sample has about 2,200 tokens, and the evaluation samples vary between 2,200 tokens to 60,000 tokens. We train for one epoch with 300 steps, use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation) and AdamW optimizer with weight decay of 0.1. We found that the optimal decimation parameters are decimation_layer = 12, $L_{base}$=2000 during training and $L_{base}$=4000 during evaluation. We intentionally decreased Lbase during training so the model could experience decimation during the training period ($L_{train}$ was a bit higher than $L_{base}$), because otherwise the training of DeciMamba and Mamba would have been identical.
@@ B.1 Comparison With Transformers
We find that vanilla Transformers of equivalent size (trained on the same dataset with a similar training recipe), have inferior length generalization abilities compared to Mamba. This is evident in all long-context tasks tested (Tables 4, 5, 6).
@@
Table 5: Comparison With Transformers - Multi-Document Retrieval. The setting is the same as in AppendixA.2. All models were trained on sequences of length 2k (10 documents). For each context length we test performance for 100 different queries and report the success rate (between 0 and 1).
Num of Docs 10 20 40 60 80 100 120 140 160 180 200 250 300
Pythia-160M 0.69
Mamba-130M 0.59 0.61 0.41 0.06
DeciMamba-130M 0.70 0.70 0.53 0.48 0.51 0.58 0.67 0.68 0.73 0.60 0.61 0.53 0.21
@@ Appendix C Hyperparameter Selection
Regarding the number of layers to decimate and the decay rate ($\beta$): The main goal of these parameters is to improve efficiency, not performance. Both parameters introduce the option to compress the sequence further by applying additional DeciMamba layers. We emphasize that this has negligible effects on performance, as shown in the sweeps below for the passkey retrieval task (Tables 7, 8). In addition, this is supported by other results, such as Multi-Document Retrieval and LongBench, which were run with this option disabled.

### v3

@@ 1 Introduction
Figure 1: Improving Mamba Extrapolation with DeciMamba. We present a novel decimation mechanism tailored for Mamba. With our method Mamba is able to process sequences that are much longer than the ones trained on while enjoying reduced inference costs. (Left) Zero-shot evaluation of an instruction-tuned 2.8B Mamba model on a subset of LongBench tasks. We show both LongBench and LongBench_e (three length groups: 0-4K, 4-8K, 8k+). MFQA, GovR, MN, TrQA, LCC, RepB, and Avg stand for MultiFieldQA, GovReport, MultiNews, TriviaQA, LCC, RepoBench-p, and Average; (Right-Top) Passkey Retrieval for Mamba-130m (Right-Bottom) Same for DeciMamba-130m. All models (130m, 2.8b) were trained on lengths of 2K tokens.
@@ 5 Experiments
Document Retrieval In this task the model receives a query and $N_{docs}$ randomly assorted documents, and its objective is to return the id of the matching document. Our data is sampled from SQuAD v2 (rajpurkar2018squadv2). We train both Mamba-130m and DeciMamba-130m to retrieve out of $N_{docs}=11$ documents (about 2k tokens), and evaluate their performance with $N_{docs}\in[11,160]$ (2k to 32k tokens). The results are presented in Table 2 (Retrieval). While the performance of Mamba and DeciMamba is similar when evaluated on sequence lengths close to those used during training, DeciMamba gains a clear advantage as the number of documents exceeds this length, achieving improved length generalization. See Appendix  A.2 for more details.
Multi-Document QA. Next, we stay in a setting similar to Sec. 1 (Document Retrieval), yet increase the level of difficulty by asking the model to answer the query in free text (instead of retrieving the id of the most relevant document). We train Mamba-130m and DeciMamba-130m on the same dataset and present the F1 score between the generated response and the ground truth answers in Tab. 2. As in the previous task, when the amount of documents is close to the amount trained on, the models perform quite similarly. When the document amount increases we can see that DeciMamba develops a clear advantage. Note that the performance of both Mamba and DeciMamba is relatively modest compared to prior work, due to the use of smaller models.
Table 2: Multi-Document Retrieval and QA. The table shows the score of Mamba and DeciMamba (+Deci) 130m models as we increase the amount of documents during evaluation. For retrieval we show the accuracy and for free-form QA we show the F1 score between the ground truth annotations and the predictions. Both models were trained using 11 documents.
# Docs: 10 20 40 60 80 100 120 140 160
Retrieval (Acc) Mamba 68.3 74.0 64.7 45.3 24.7 5.7 1.0 0.3 0.3
+Deci 67.7 77.3 68.7 65.3 49.7 37.0 26.3 16.7 5.3
QA (F1) Mamba 25.3 22.8 23.6 25.8 19.3 7.5 4.1 1.1 0.6
+Deci 26.2 24.4 25.3 22.74 19.1 18.1 12.9 10.0 5.0
@@ A.1 Passkey Retrieval
A.2 Multi-Document Retrieval and QA
We train each model with data from SQuAD v2 (rajpurkar2018squadv2), which provides examples in the form of (Query, Document, Answer). Our training samples have the following form: $N_{docs}\times$$<$$Document$$>$; $<$$Answer$$>$, where $<$$Document$$>$ can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. $<$$Answer$$>$ holds the id of the golden document. In our setting $N_{docs}=11$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During Evaluation we use the same setting but vary the value of $N_{docs}$, between 11 and 160. We note that the length of an average document in SQuAD is a bit smaller than 200 tokens, so our average training sample has about 2,200 tokens, and the evaluation samples vary between 2,200 tokens to 32,000 tokens. We train for two epochs (1500 steps in each), use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation), and AdamW optimizer with weight decay of 0.1. We found that the optimal decimation parameters are decimation_layer = 12, $L_{base}$=2000 during training and $L_{base}$=4000 during evaluation. We intentionally decreased Lbase during training so the model could experience decimation during the training period ($L_{train}$ was a bit higher than $L_{base}$), because otherwise the training of DeciMamba and Mamba would have been identical. The displayed results are the average performance over 3 different training seeds.
@@ B.1 Comparison With Transformers
We find that vanilla Transformers of equivalent size (trained on the same dataset with a similar training recipe), have inferior length generalization abilities compared to Mamba. This is evident in multiple long-context tasks (Tables 4, 5).
@@ Appendix C Hyperparameter Selection
Regarding the number of layers to decimate and the decay rate ($\beta$): The main goal of these parameters is to improve efficiency, not performance. Both parameters introduce the option to compress the sequence further by applying additional DeciMamba layers. We emphasize that this has negligible effects on performance, as shown in the sweeps below for the passkey retrieval task (Tables 6, 7). In addition, this is supported by other results, such as Multi-Document Retrieval and LongBench, which were run with this option disabled.